OpenStudy (anonymous):

find the domain of sq root x / sq root x^2-6x+8

OpenStudy (anonymous):

\[\sqrt{x}\div \sqrt{x^2-6x+8} what is the domain\]

OpenStudy (pradius):

\[x ^{2}-6x+8 \ge 0\]